[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Aug 31, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Thu Aug 31 12:31:59 UTC 2017


Here is today's summary of NHC computer operations:
--- Pablo

----------

TAFB:
1.) From Morning morning Round Ups. 0345Z Rainfall product did not run for  
TS Harvey. 0345Z Rainfall product for Pot. TC 10 is named "Tropical Storm  
10" in the product. Hugh inquires if we looked into this. Matt Sardi, did  
you have a chance to do so? I know you guys were busy restoring data flow  
into AWIPS yesterday.

2) Still no nwps_ep or nwps_al into NAWIPS. Not sure the status of this.  
Again, this is important in calculating the 12 ft sea radii.

I am asking Matt Sardi to look into this as Chris Mello is tied up working  
with Monica on a long list of overdue items. That is why Monica is here  
this week as well.

Meanwhile, please know you can calculate 12 ft seas radii using the AWIPS  
D2D Distance and/or home tools looking at NWPS output on D2D. I will be  
happy to show you while the guys have time to restore this in NAWIPS as  
well. Jeff and Eric can show you as well. It would not be a bad idea to  
start making the transition now anyhow as NAWIPS is faced out over the next  
few years.

I ask everyone to be somewhat patient about this. We have 8 standing RFCs,  
multiple priority tickets including replacement of workstations, monitors,  
and graphic cards upgrades over the coming weeks that should all address a  
lot of issues we keep seeing. You will be seeing a lot of changes over the  
coming weeks but activity and CWD declarations are holding us back.
HSU:
1.) Tuna is VERY slow. Brian M rebooted but no effect. Doug G updated  
drives and workstation is running much better.

2.) Could someone install the same top monitor that the Atlantic desk has  
at the ridges workstation on sawfish? Yes we are working on this. But CWD  
has a lot of this scheduled work on hold. Please video card upgrades needed  
in support of new higher res monitors and higher resolution data being  
displayed. See comments above under TAFB section.

3.) SHIPS/LGEM did not initially run for Irma. Nothing printed out from the  
GUI, and files online were blank. Also not in the a-deck. Guidance was  
originally run on the backup ATCF. Richard reran the guidance on the  
Atlantic ATCF, and SHIPS/LGEM became available. Craig Mattock will follow  
this up.
Other items:
1.) We had glitches again with security system yesterday but William reset  
it. We are still awaiting a brand new higher power switch to be delivered  
by the end of the week.
NHC RFCs:
See list of pending RFCs sent with Monday's Morning Rounds.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170831/8ec2580a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list